OpenStudy (anonymous):

a varies directly as b and inversely as the square of c, and a=10 when b=8 and c=2. Find a when b=20 and c=5.

OpenStudy (anonymous):

the equation for directly is x=ky so for this it would be a=bk and for inverse it is x=y/k so it would be a=c/k you plug it in from there

OpenStudy (anonymous):

But the inverse formula is y=k/x ....

OpenStudy (anonymous):

a = kb + k/c^2 10 = 8k + k / 4 40 = 32k + k k = 40/33 a = (40/33)b + (40/33)/ c^2 when b=20 and c = 5 a = (40/33)*20 + (40/33) / 25 = 800/33 + 40 / (33*25)

OpenStudy (anonymous):

- i get that to 24.29